Using the Density of water of 1 g/cm3 and Given the following dimensions, calculate the mass of the water in grams for water in the Fish tank. 15 inches high x .075 m wide x 30 cm long. (I am really lost can you show me how you got to your answer too?)
Answer by Alan3354(69443)

Using the Density of water of 1 g/cm3 and Given the following dimensions, calculate the mass of the water in grams for water in the Fish tank. 15 inches high x .075 m wide x 30 cm long.
-------
Change all dimensions to cm
15 in * 2.54 cm/in = 38.1 cm
0.75 meters * 100 cm/m = 75 cm
-------
Vol = L*W*H = 38.1*75*30 = 85725 cc
85725 cc * 1 gm/cc = 85725 grams